The frequency of violet light is 7.31 x 1014 s-1 and that of red light is 4.57 x 1014 s-1.
(a) Calculate the wavelength of each color.
(b) Calculate the energy, in Joules, of an individual photon in each of the colors of light.
Answer #1

a) wavelength = Speed of light/frequency

Wavelength 1 = 3×108/7.31×1014 = 4.1×10-7 m

Wavelength 2 = 3×108/4.57×1014 = 6.56×10-7 m

b) E = hv

E1 = 6.626×10-34×7.31×1014 = 4.84×10-21 J

E2 = 6.626×10-34×4.57×1014 = 3.03×10-21 J

Please give feedback!